OpenStudy (anonymous):

What is the equation of the line which includes points (1, 5) and (-2, -7)? A. y = -4x - 1 B. y equals negative four x plus one half C. y = 4x + 1 D. y equals negative one fourth x plus one

OpenStudy (anonymous):

calculate the slope between these two points... wat u got?

OpenStudy (anonymous):

\(\large slope=m=\frac{y_2-y_1}{x_2-x_1}=\frac{5-(-7)}{1-(-2)}= \) ????

OpenStudy (anonymous):

4